3×1.6^4+9×1.6×2/6-18.4
Answer by fractalier(6550)

Using PEMDAS (order of operations) rules
3 × 1.6^4 + 9×1.6×2/6 - 18.4 =
3 x 6.5536 + 28.8/6 - 18.4 =
19.6608 + 4.8 - 18.4 =
6.0608